# The Nondecreasing Rank

### v1

@@ 7.1 The Maximum ND Rank
Assume that ${\mathbf{T}}=\sum_{i=1}^{r}{\mathbf{a}}_{i}\otimes{\mathbf{b}}_{i}$ is a minimum ND rank decomposition, with ${\mathbf{a}}_{i}=\sum_{j=1}^{2^{p-1}}\lambda_{ij}{\mathbf{v}}_{j}$ and ${\mathbf{b}}_{i}=\sum_{l=1}^{2^{p-1}}\mu_{il}{\mathbf{v}}_{l}$, $\lambda_{ij},\mu_{il}\geq 0$. For any pair of extremal rays ${\mathbf{v}}_{l}\neq{\mathbf{v}}_{j}$ choose a facet $F$ with corresponding dual vector $\mathbf{h}$} so that ${\mathbf{v}}_{l}\in F$ but ${\mathbf{v}}_{j}\notin F$. Let $\mathbf{h}$$^{\prime}$ be the dual vector corresponding to the facet $F^{\prime}$ opposite of $F$. We compute
$0={\mathbf{h}}^{\intercal}\sum_{i=1}^{2^{p-1}}({\mathbf{v}}_{i}\otimes{\mathbf{v}}_{i}){\mathbf{h}}^{\prime}={\mathbf{h}}^{\intercal}\big{(}\sum_{i=1}^{r}\sum_{j=1}^{2^{p-1}}\sum_{l=1}^{2^{p-1}}\lambda_{ij}\mu_{il}({\mathbf{v}}_{j}\otimes{\mathbf{v}}_{l})\big{)}{\mathbf{h}}^{\prime}\geq\sum_{i=1}^{r}\lambda_{ij}\mu_{il}({\mathbf{h}}^{\intercal}{\mathbf{v}}_{j})({\mathbf{v}}_{l}^{\intercal}{\mathbf{h}}^{\prime}).$
@@
The first equality follows because every extremal ray in $\mathcal{C}(\mathcal{P})$ must either be in the facet $F$ or $F^{\prime}$. By construction $({\mathbf{h}}^{\intercal}{\mathbf{v}}_{j})({\mathbf{v}}_{l}^{\intercal}{\mathbf{h}}^{\prime})>0$, which implies that $\lambda_{ij}\mu_{il}=0$ whenever $j\neq l$. We conclude that ${\mathbf{a}}_{i}\otimes{\mathbf{b}}_{i}\,\propto\,{\mathbf{v}}_{j_{i}}\otimes{\mathbf{v}}_{j_{i}}$ for some index $j_{i}$, with ${\mathbf{T}}=\sum_{i=1}^{r}c_{i}{\mathbf{v}}_{j_{i}}\otimes{\mathbf{v}}_{j_{i}}$. By contradiction assume that the extremal ray ${\mathbf{v}}_{l}$ does not appear in the set $\{{\mathbf{v}}_{j_{i}}:i=1,\ldots,r\}$. Take $F$ to be a facet that does not contain ${\mathbf{v}}_{l}$, where $\mathbf{h}$} is the dual vector corresponding to its supporting hyperplane, and $F^{\prime}$ is the opposite facet. Then
$\sum_{i:{\mathbf{v}}_{i}\in F^{\prime}}{\mathbf{v}}_{i}=\sum_{i=1}^{2^{p-1}}({\mathbf{v}}_{i}\otimes{\mathbf{v}}_{i})({\mathbf{h}})=\sum_{i=1}^{r}c_{i}({\mathbf{v}}_{j_{i}}\otimes{\mathbf{v}}_{j_{i}})({\mathbf{h}})=\sum_{i:{\mathbf{v}}_{j_{i}}\in F^{\prime}}c_{i}{\mathbf{v}}_{i}.$
The vector $\sum_{i:{\mathbf{v}}_{i}\in F^{\prime}}{\mathbf{v}}_{i}$ is in the relative interior of $F^{\prime}$ while $\sum_{i:{\mathbf{v}}_{j_{i}}\in F^{\prime}}c_{i}{\mathbf{v}}_{i}$ is not, a contradiction. ∎
We remark that both Theorems 6 and 7 show that the upper bound in Lemma 6 is attained. It is an open question as to whether this bound is attained in the matrix setting for every order cone and for more general polyhedral cones.
@@ 7.2 A Matrix Tri-Factorization Formulation
A succinct representation of a rank-$r$, nonnegative matrix factorization of ${\mathbf{T}}\in\mathbb{R}^{p_{1}\times p_{2}}_{+}$ is ${\mathbf{T}}={\mathbf{A}}_{1}{\mathbf{A}}_{2}^{\intercal}$ where ${\mathbf{A}}_{j}\in\mathbb{R}_{+}^{p_{j}\times r}$. If ${\mathbf{T}}=\sum_{i=1}^{r}{\mathbf{a}}^{(i)}_{1}\otimes{\mathbf{a}}_{2}^{(i)}$ is an ND factorization we can likewise write this as ${\mathbf{T}}={\mathbf{A}}_{1}{\mathbf{A}}_{2}^{\intercal}$ where ${\mathbf{A}}_{j}\in\mathbb{R}^{p_{j}\times r}$ has columns that are given by the ${\mathbf{a}}_{j}^{(i)}$s for $i=1,\ldots,p_{j}$. Let ${\mathbf{V}}_{j}\in\mathbb{R}^{p_{i}\times q_{j}}$ be a matrix that has columns that are equal to the $q_{j}$ extremal rays in $\mathcal{C}(\mathcal{P}_{j})$. As each ${\mathbf{a}}^{(i)}_{j}$ is in the conical hull of the columns of ${\mathbf{V}}_{j}$ there exist nonnegative matrices ${\mathbf{H}}_{i}\in\mathbb{R}_{+}^{q_{i}\times r}$ where ${\mathbf{A}}_{j}={\mathbf{V}}_{j}{\mathbf{H}}_{j}$ for $j=1,2$. Defining ${\mathbf{H}}={\mathbf{H}}_{1}{\mathbf{H}}_{2}^{\intercal}\in\mathbb{R}^{q_{1}\times q_{2}}_{+}$ the ND factorization can be written as
${\mathbf{T}}={\mathbf{V}}_{1}{\mathbf{H}}_{1}{\mathbf{H}}_{2}^{\intercal}{\mathbf{V}}_{2}^{\intercal}={\mathbf{V}}_{1}{\mathbf{H}}{\mathbf{V}}_{2}^{\intercal}.$
@@
The nondecreasing rank of a matrix ${\mathbf{T}}\in\mathcal{N}_{<\infty}$ is the smallest nonnegative rank of a matrix ${\mathbf{H}}\in\mathbb{R}^{q_{1}\times q_{2}}$ that satisfies the equation ${\mathbf{T}}={\mathbf{V}}_{1}{\mathbf{H}}{\mathbf{V}}_{2}^{\intercal}$, where ${\mathbf{V}}_{j}\in\mathbb{R}^{p_{j}\times q_{j}}$ are fixed matrices with rows that are equal to the $q_{j}$ extremal rays of $\mathcal{C}(\mathcal{P}_{j})$.
@@
$\begin{bmatrix}1&1&2\\ 1&2&2\\ 2&2&4\end{bmatrix}=\begin{bmatrix}0&1&0&1\\ 0&0&1&1\\ 1&1&1&1\end{bmatrix}\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\end{bmatrix}\begin{bmatrix}0&0&1\\ 1&0&1\\ 0&1&1\\ 1&1&1\end{bmatrix}={\mathbf{V}}_{1}{\mathbf{H}}{\mathbf{V}}_{2}^{\intercal},$
where $\mathbf{H}$} has a nonnegative rank of four and the rows of ${\mathbf{V}}_{1}={\mathbf{V}}_{2}$ are equal to the four extremal rays of the collider $\mathcal{C}(\mathcal{P}_{j})$ outlined in Lemma 2.
@@ 7.3 Typical ND Ranks
A typical nondecreasing rank is defined as a number $r$ such that $\mathcal{N}_{r}$ has a non-empty interior. The probabilistic interpretation of a typical rank is that if a matrix $\mathbf{X}$} is drawn from a distribution supported on $\mathcal{N}_{<\infty}$ that has a density with respect to the Lebesgue measure then $\text{Pr}(\text{NDrank}({\mathbf{X}})=r)>0$ when $r$ is a typical rank. In this section we determine the typical ranks for the two matrix settings outlined in Theorems 6 and 7, as well as for any other setting where the maximum ND rank is known. Unlike the typical nonnegative rank, there can be multiple, typical nondecreasing ranks when $k=2$, a consequence of the maximum ND rank potentially being larger than $\min(p_{1},p_{2})$. We begin by extending two results on typical nonnegative ranks to the nondecreasing rank. The real, typical rank in $\otimes_{j=1}^{k}\mathbb{R}^{p_{j}}$ refers to any $r$ where the set of real, rank-$r$ tensors in $\otimes_{j=1}^{k}\mathbb{R}^{p_{j}}$ has a non-empty interior.
Theorem 8 (Theorem 2 [10]).
The minimum, typical, ND rank in $\otimes_{j=1}^{k}\mathcal{C}(\mathcal{P}_{j})$ is equal to the minimum, typical, real rank in $\otimes_{j=1}^{k}\mathbb{R}^{p_{j}}$.
Proof.
The proof of this result in [10] relies on the fact that the set of tensors of rank at most $r$ is a semialgebraic set. A semialgebraic set in $\mathbb{R}^{p}$ is comprised of finite unions, intersections, and complements of sets of the form $\{{\mathbf{x}}:f({\mathbf{x}})\leq 0\}$, where $f$ is a multivariate polynomial. The set $\mathcal{N}_{\leq r}$ is semialgebraic by the Tarski-Seidenberg theorem [11, Thm 2.2.7] as it is the image of the semialgebraic set $(\mathcal{C}(\mathcal{P}_{1})\oplus\cdots\oplus\mathcal{C}(\mathcal{P}_{k}))^{r}\subseteq(\mathbb{R}^{p_{1}}\oplus\cdots\oplus\mathbb{R}^{p_{k}})^{r}$ under the polynomial map $({\mathbf{v}}^{(11)},\ldots,{\mathbf{v}}^{(rk)})\mapsto\sum_{i=1}^{r}\otimes_{j=1}^{k}{\mathbf{v}}^{(ij)}$. ∎
For $p_{1}\times p_{2}$ matrices the only typical real rank is $\min(p_{1},p_{2})$, which by the above theorem is always a typical ND rank.
Theorem 9 (Theorem 2.2 [8]).
If $r$ is a typical ND rank less than the maximum, typical, ND rank then $r+1$ is also a typical ND rank.
Proof.
The proof provided in [8] also only relies on the fact that $\mathcal{N}_{\leq r}$ is semialgebraic. ∎
From Theorem 9 it remains to find the largest typical ND rank to completely determine all possible typical ranks.
@@
Let $m$ be the maximum ND rank and choose a $\mathbf{T}$} with $\text{NDrank}({\mathbf{T}})=m$. If $B_{n^{-1}}({\mathbf{T}})$ is a ball of radius $n^{-1}$ centered at $\mathbf{T}$}, and if the maximum ND rank was not a typical rank, the set $B_{n^{-1}}({\mathbf{T}})\cap\mathcal{N}_{<\infty}$, that has a non-empty interior, would contain an ${\mathbf{S}}_{n}\in\mathcal{N}_{\leq m-1}$. Taking $n\rightarrow\infty$ gives a sequence in $\mathcal{N}_{\leq m-1}$ with ${\mathbf{S}}_{n}\rightarrow{\mathbf{T}}$. As the set $\mathcal{N}_{\leq m-1}$ is closed (see Theorem 13) this is a contradiction to ${\mathbf{T}}\in\mathcal{N}_{m}$. ∎
@@
If $\mathcal{P}=\{x_{1},\ldots,x_{p}\}$ where $x_{i}\prec x_{p}$ for all $i<p$ then $r$ is a typical rank in $\mathcal{C}(\mathcal{P})\otimes\mathcal{C}(\mathcal{P})$ if and only if $p\leq r\leq 2^{p-1}$.

### v2

@@ 7.1 The Maximum ND Rank
Assume that ${\mathbf{T}}=\sum_{i=1}^{r}{\mathbf{a}}_{i}\otimes{\mathbf{b}}_{i}$ is a minimum ND rank decomposition, with ${\mathbf{a}}_{i}=\sum_{j=1}^{2^{p-1}}\lambda_{ij}{\mathbf{v}}_{j}$ and ${\mathbf{b}}_{i}=\sum_{l=1}^{2^{p-1}}\mu_{il}{\mathbf{v}}_{l}$, $\lambda_{ij},\mu_{il}\geq 0$. For any pair of extremal rays ${\mathbf{v}}_{l}\neq{\mathbf{v}}_{j}$ choose a facet $G$ with corresponding dual vector $\mathbf{h}$} so that ${\mathbf{v}}_{l}\in G$ but ${\mathbf{v}}_{j}\notin G$. Let $\mathbf{h}$$^{\prime}$ be the dual vector corresponding to the facet $G^{\prime}$ opposite of $G$. We compute
$0={\mathbf{h}}^{\intercal}\sum_{i\in I}\alpha_{i}({\mathbf{v}}_{i}\otimes{\mathbf{v}}_{i}){\mathbf{h}}^{\prime}={\mathbf{h}}^{\intercal}\big(\sum_{i=1}^{r}\sum_{j=1}^{2^{p-1}}\sum_{l=1}^{2^{p-1}}\lambda_{ij}\mu_{il}({\mathbf{v}}_{j}\otimes{\mathbf{v}}_{l})\big){\mathbf{h}}^{\prime}\geq\sum_{i=1}^{r}\lambda_{ij}\mu_{il}({\mathbf{h}}^{\intercal}{\mathbf{v}}_{j})({\mathbf{v}}_{l}^{\intercal}{\mathbf{h}}^{\prime}).$
@@
The first equality follows because every extremal ray in $\mathcal{C}(\mathcal{P})$ must be in only one of the facets $G$ or $G^{\prime}$. The inequality is a result of the middle expression being a sum of nonnegative terms. By construction $({\mathbf{h}}^{\intercal}{\mathbf{v}}_{j})({\mathbf{v}}_{l}^{\intercal}{\mathbf{h}}^{\prime})>0$, which implies that $\lambda_{ij}\mu_{il}=0$ whenever $j\neq l$. We conclude that ${\mathbf{a}}_{i}\otimes{\mathbf{b}}_{i}\,\propto\,{\mathbf{v}}_{j_{i}}\otimes{\mathbf{v}}_{j_{i}}$ for some index $j_{i}$, with ${\mathbf{T}}=\sum_{i=1}^{r}c_{i}{\mathbf{v}}_{j_{i}}\otimes{\mathbf{v}}_{j_{i}}$ and $c_{i}>0$. Let $\mathbf{h}$} be a dual vector corresponding to the hyperplane that supports the facet $F^{\prime}$. Then
$\sum_{i\in I_{F}}\alpha_{i}{\mathbf{v}}_{i}=\sum_{i\in I}\alpha_{i}({\mathbf{v}}_{i}\otimes{\mathbf{v}}_{i})({\mathbf{h}})=\sum_{i=1}^{r}c_{i}({\mathbf{v}}_{j_{i}}\otimes{\mathbf{v}}_{j_{i}})({\mathbf{h}})=\sum_{i:{\mathbf{v}}_{j_{i}}\in F}c_{i}{\mathbf{v}}_{j_{i}},$
where without loss of generality it is assumed that $\mathbf{h}$} is scaled so that ${\mathbf{h}}^{\intercal}{\mathbf{v}}_{i}=1$ for ${\mathbf{v}}_{i}\in F$. By how the vector $\sum_{i\in I_{F}}\alpha_{i}{\mathbf{v}}_{i}$ was constructed, there must exist at least p1p-1 terms in the sum $\sum_{i:{\mathbf{v}}_{j_{i}}\in F}c_{i}{\mathbf{v}}_{j_{i}}$. A symmetric argument shows that the set of indices $\{i:{\mathbf{v}}_{j_{i}}\in F^{\prime}\}$ has a size of at least p1p-1. As the facets $F$ and $F^{\prime}$ do not share any extremal rays, we conclude that the sum $\sum_{i=1}^{r}c_{i}{\mathbf{v}}_{j_{i}}\otimes{\mathbf{v}}_{j_{i}}$ must include at least 2(p1)2(p-1) terms and so $r\geq 2(p-1)$.
In the case when $p=3$, Lemma 6 shows that the maximum rank is also at most 4=2(31)4=2(3-1), which is the number of vertices in a two-dimensional cube. ∎
The problem of finding the maximum rank of tensor decompositions for arbitrary order cones and for more general polyhedral cones remains an open question.
@@ 7.2 A Matrix Tri-Factorization Formulation
A succinct representation of a rank-$r$, nonnegative matrix factorization of ${\mathbf{T}}\in\mathbb{R}^{p_{1}\times p_{2}}_{+}$ is ${\mathbf{T}}={\mathbf{A}}^{(1)}({\mathbf{A}}^{(2)})^{\intercal}$ where ${\mathbf{A}}^{(j)}\in\mathbb{R}_{+}^{p_{j}\times r}$. If ${\mathbf{T}}=\sum_{i=1}^{r}{\mathbf{a}}^{(i1)}\otimes{\mathbf{a}}_{2}^{(i2)}$ is an ND factorization we can likewise write this as ${\mathbf{T}}={\mathbf{A}}^{(1)}({\mathbf{A}}^{(2)})^{\intercal}$ where ${\mathbf{A}}^{(j)}\in\mathbb{R}^{p_{j}\times r}$ has columns that are given by the ${\mathbf{a}}^{(ij)}$s for $i=1,\ldots,r$. Let ${\mathbf{V}}^{(j)}\in\mathbb{R}^{p_{i}\times q_{j}}$ be a matrix that has columns that are equal to the $q_{j}$ extremal rays in $\mathcal{C}(\mathcal{P}_{j})$. As each ${\mathbf{a}}^{(ij)}$ is in the conical hull of the columns of ${\mathbf{V}}^{(j)}$ there exist nonnegative matrices ${\mathbf{H}}^{(i)}\in\mathbb{R}_{+}^{q_{i}\times r}$ where ${\mathbf{A}}^{(j)}={\mathbf{V}}^{(j)}{\mathbf{H}}^{(j)}$ for $j=1,2$. Defining ${\mathbf{H}}={\mathbf{H}}^{(1)}({\mathbf{H}}^{(2)})^{\intercal}\in\mathbb{R}^{q_{1}\times q_{2}}_{+}$ the ND factorization can be written as
${\mathbf{T}}={\mathbf{V}}^{(1)}{\mathbf{H}}^{(1)}({\mathbf{H}}^{(2)})^{\intercal}({\mathbf{V}}^{(2)})^{\intercal}={\mathbf{V}}^{(1)}{\mathbf{H}}({\mathbf{V}}^{(2)})^{\intercal}.$
@@
The nondecreasing rank of a matrix ${\mathbf{T}}\in\mathcal{N}_{<\infty}$ is the smallest nonnegative rank of a matrix ${\mathbf{H}}\in\mathbb{R}^{q_{1}\times q_{2}}$ that satisfies the equation ${\mathbf{T}}={\mathbf{V}}^{(1)}{\mathbf{H}}({\mathbf{V}}^{(2)})^{\intercal}$, where ${\mathbf{V}}^{(j)}\in\mathbb{R}^{p_{j}\times q_{j}}$ are fixed matrices with columns that are equal to the $q_{j}$ extremal rays of $\mathcal{C}(\mathcal{P}_{j})$.
@@
$\begin{bmatrix}2&1&2\\ 1&2&2\\ 2&2&4\end{bmatrix}=\begin{bmatrix}0&1&0&1\\ 0&0&1&1\\ 1&1&1&1\end{bmatrix}\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\end{bmatrix}\begin{bmatrix}0&0&1\\ 1&0&1\\ 0&1&1\\ 1&1&1\end{bmatrix}={\mathbf{V}}^{(1)}{\mathbf{H}}({\mathbf{V}}^{(2)})^{\intercal},$
where $\mathbf{H}$} has a nonnegative rank of four and the rows of ${\mathbf{V}}^{(1)}={\mathbf{V}}^{(2)}$ are equal to the four extremal rays of the collider $\mathcal{C}(\mathcal{P}_{j})$ outlined in Lemma 2.
@@ 7.3 Typical ND Ranks
A typical nondecreasing rank is defined as a number $r$ such that $\mathcal{N}_{r}$ has a non-empty interior. The probabilistic interpretation of an typical ND rank is that if a matrix $\mathbf{X}$} is drawn from a distribution supported on $\mathcal{N}_{<\infty}$ that has a density with respect to the Lebesgue measure then $\text{Pr}(\text{NDrank}({\mathbf{X}})=r)>0$ when $r$ is a typical rank. In this section we determine the typical ranks for the two matrix settings outlined in Theorems 6 and 7, as well as for any other setting where the maximum ND rank is known.
@@
Let $m$ be the maximum ND rank and choose a $\mathbf{T}$} with $\text{NDrank}({\mathbf{T}})=m$. Let $B_{n^{-1}}({\mathbf{T}})$ be an open ball of radius $n^{-1}$ centered at $\mathbf{T}$}. The set $B_{n^{-1}}({\mathbf{T}})\cap\mathcal{N}_{<\infty}$ has a non-empty interior. If ${\mathbf{T}}\in\text{int}(\mathcal{N}_{<\infty})$ this is immediate. Otherwise, if $\mathbf{T}$} is in the boundary of $\mathcal{N}_{<\infty}$ take an ${\mathbf{R}}\in\text{int}(\mathcal{N}_{<\infty})$, which exists because $\mathcal{N}_{<\infty}$ is proper, and consider the line segment $\lambda{\mathbf{T}}+(1-\lambda){\mathbf{R}}$, $\lambda\in[0,1]$. For all $\lambda\in[0,1)$ the point $\lambda{\mathbf{T}}+(1-\lambda){\mathbf{R}}$ is in $\text{int}(\mathcal{N}_{<\infty})$, and for large enough $\lambda$ this point is also in $B_{n^{-1}}({\mathbf{T}})$. this provides a point in $B_{n^{-1}}({\mathbf{T}})\cap\text{int}(\mathcal{N}_{<\infty})$. Now if the maximum ND rank was not a typical rank, the set $B_{n^{-1}}({\mathbf{T}})\cap\mathcal{N}_{<\infty}$, that has a non-empty interior, would contain an ${\mathbf{S}}_{n}\in\mathcal{N}_{\leq m-1}$. Taking $n\rightarrow\infty$ gives a sequence in $\mathcal{N}_{\leq m-1}$ with ${\mathbf{S}}_{n}\rightarrow{\mathbf{T}}$. As the set $\mathcal{N}_{\leq m-1}$ is closed (see Theorem 12) this is a contradiction to ${\mathbf{T}}\in\mathcal{N}_{m}$. Note that this proof also shows that the set $\mathcal{N}_{m}$ is an open set in the subspace topology of $\mathcal{N}_{<\infty}$ since there must exist an $n$ with $B_{n^{-1}}({\mathbf{T}})\cap\mathcal{N}_{<\infty}\subset\mathcal{N}_{m}$. ∎
The next theorem extends results that have appeared in [43, 54] for the nonnegative rank and completely characterizes all possible typical ND ranks. A typical real rank in $\otimes_{j=1}^{k}\mathbb{R}^{p_{j}}$ refers to any $r$ where the set of real, rank-$r$ tensors in $\otimes_{j=1}^{k}\mathbb{R}^{p_{j}}$ has a non-empty interior.
Theorem 9.
Let $r_{Re}$ be the minimum, typical, real rank for tensors in $\mathbb{R}^{p_{1}\times\cdots\times p_{k}}$. Any $r$ with $r_{Re}\leq r\leq\mathrm{maxNDrank}$ is a typical ND rank.
Proof.
We adapt the proof provided in Theorem 3.8 of [54] by replacing all instances of the the cone of nonnegative vectors $\mathbb{R}^{p_{j}}_{+}$ with the order cone $\mathcal{C}(\mathcal{P}_{j})$. This proof does not rely on any of the structure of the cone $\mathbb{R}^{p_{j}}_{+}$; the only place in this proof where the cone comes into play is in showing that the maximum rank is a typical nonnegative rank. In the ND rank case the analogous result is shown in Theorem 8. Theorem 3.8 also depends on Lemmas 3.4-3.7 in [54]. Similarly, none of these lemmas rely on the specific structure of $\mathbb{R}_{+}^{p_{i}}$ and also apply to $\mathcal{C}(\mathcal{P}_{j})$ and the ND rank. ∎
When considering rank-$r$ factorizations over the complex numbers there is only one typical rank, also referred to as the generic rank [10]. In general, there can be more than one typical, real or nonnegative rank [53]. However, for matrices, since the maximum rank of a $p_{1}\times p_{2}$ matrix is equal the minimum typical rank of $\min(p_{1},p_{2})$, this is the only typical, real or nonnegative rank. As the maximum ND rank can be larger than $\min(p_{1},p_{2})$ there can be more than one typical ND rank in the matrix setting, as is illustrated by the following corollaries.
@@
Proof.
This is a direct consequence of Theorems 6, 8, and 9; collectively these results show that any rank greater than or equal to $\min(p_{1},p_{2})$ and less than or equal to the maximum rank of $\min(q_{1},q_{2})$ is a typical rank. ∎
@@
If $\mathcal{P}=\{x_{1},\ldots,x_{p}\}$ where $x_{i}\prec x_{p}$ for all $i<p$ then $r$ is a typical rank in $\mathcal{C}(\mathcal{P})\otimes\mathcal{C}(\mathcal{P})$ if $p\leq r\leq 2(p-1)$.
Proof.
The proof is the same as in Corollary 1 except that the maximum rank is now bounded below by 2(p1)2(p-1) in Theorem 7. ∎